OpenStudy (anonymous):

f(x,y)=y/x^2 a.) find functions domain and range b.) describe the functions level curves

OpenStudy (dan815):

x=/=0

OpenStudy (dan815):

level curves means what does that graph look like at different heights which is basically different planes of Z

OpenStudy (dan815):

at f(x,y)=Z=1=y/x^2 when you got z=1 u have y=x^2 z=2 u have y=2*x^2 Z=3...........y=3*x^2

OpenStudy (anonymous):

@dan815 oh i see! however not sure i understand what the domain and range would be?